OpenStudy (anonymous):

find f^-1(x): f(x)=x^3+1

OpenStudy (anonymous):

f(x) = x³ + 1 y = x³ + 1 x = y³ + 1 x - 1 = y³ y = ³√(x - 1) \[f ^{-1}(x) = \sqrt[3]{x-1}\]